find the present and future value of 1000 received every month end for 20 years if the interest rate is j12= 12% per annum
1
Expert's answer
2020-06-18T18:28:40-0400

To Calculate Present Value:



"PV = CF\left[\dfrac{1 - (1 + i)^{-n}}{i}\right]"

In our question, we have:


"CF = 1,000\\[0.3cm]
i = \dfrac {0.12}{12} = 0.01\\[0.3cm]
n = 20\times 12 = 240"


Therefore:



"PV = 1,000\left[\dfrac{1 - (1 + 0.01)^{-240}}{0.01}\right]\\[0.3cm]
\color{red}{PV = 90,819.42}"